A gas has a volume of 4.10 L at 0 ∘C. What final temperature, in degrees Celsius, is needed to change the volume of the gas to each of the following, if n and P do not change?

1.00 L Express your answer using three significant figures.

1800 mL Express your answer using two significant figures

10.5 L Express your answer using three significant figures.

54.0 mL Express your answer using three significant figures.
Answer #1

we will use charts low, which states that At constant pressure, the volume of gas is directly propostional to Kenin temperatu☺ r₂ = 10.5L 2866.5k.L T2 = 273k x10.52 4.10L -= 4.10L = 699.14K : 699.14 K – 273 = 1426°C/ 4) Van 54 m2 = 54/1000 =0.05.4 L